OpenStudy (anonymous):

Solve the system by substitution. -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

Lets take 2x + 2z = 4 Dividing each term by 2 we get x + z = 2 => x = 2 - z

OpenStudy (anonymous):

Now lets substitute x in any one of the first two equations. Lets take -x -y -z = -8 Multiply each term with minus sign we get x + y + z = 8 now get x = 2-z in place of x in the above equation

OpenStudy (anonymous):

2 - z + y + z = 8 => 2 + y = 8 => y = 6

OpenStudy (anonymous):

Now substitute x and y in second equation that is -4x + 4y + 5z = 7 -4(2 - z) + 24 + 5z = 7 -8 + 4z + 24 + 5z = 7 9z = 7 - 16 9z = -9 z = -1

OpenStudy (anonymous):

Now Substitute z = -1 in x = 2 - z we get x = 2 - (-1) = 2 + 1 = 3

OpenStudy (anonymous):

Thank you so much I was stuck on this question for the longest. Can you help me with another one?

OpenStudy (anonymous):

Sure np

OpenStudy (anonymous):

Solve the system of elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

Lets take first and 3rd equations and add them

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok did you multiply the second equation?

OpenStudy (anonymous):

No I didnt I just added first and 3rd equations

OpenStudy (anonymous):

Now lets subtract first two equations

OpenStudy (anonymous):

-2x + 2y + 3z = 0 -2x -y + z = -3 (+) (+) (-) (+) Subtracting these 2 3y +2z = 3

OpenStudy (anonymous):

Now take equation 4 and 5 which we got as resultant in above two equations

OpenStudy (anonymous):

5y + 6z = 5 3y + 2z = 3 Now Multiply equation 3y + 2z = 3 with 2 and subtract from 5y + 6z = 5 5y + 6z = 5 9y + 6z = 9 (-) (-) (-) ............................ -4y = -4 => y = 1 Substitute y = 1 in 3y + 2z = 3 3 + 2z = 3 2z = 0 => z = 0

OpenStudy (anonymous):

Now Substitute y and z values in any of the given 3 equations to get x value

OpenStudy (anonymous):

-2x + 2y + 3z = 0 -2x + 2 =0 -2x = -2 x = 1

OpenStudy (anonymous):

1 ,1 , 0 are the solutions for x, y , z

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

Anytime !